# StateAct: State Tracking and Reasoning for Acting and Planning with Large Language Models

### v1

@@ 4.4 Models
In line with previous work we focus our main experiments are based on the API based LLMs to compare performance fairly. Concretely, ‘gpt-3.5’ level models were used in previous work including current state-of-the-art ActRe (Yang et al., 2024b) and state-of-the-art for in-context learning ReAct (Yao et al., 2023b). Many OpenAI models have become deprecated. Notably, all models from ReAct and AdaPlannerSun et al. (2023) davinci-002, gpt-3.5-turbo-0301 and gpt-3.5-turbo-0613 are deprecated now. Therefore, we re-implemented ReAct and ran the experiments to determine the most suitable model, see Table 1. We establish that gpt-3.5-turbo-1106 is the best performing (from the ones that remain available) on ReAct and we therefore chose this one. In our work, we additionally compare our method using newer state-of-the-art models to show that our method generalises to new model. Specificially, we use the gpt-4o-mini and Mixtral(Jiang et al., 2024) model.. We note that full GPT-4 level models are prohibitively expensive [^4] A single evaluation run on alfworld costs approx. \$8 using gpt-3.5 and ReAct, gpt-4 would cost 10+ times more.. We use temperature 0 for all experiments and sample only the top 1 response, see Appendix C the exact settings.
@@
Table 1: ReAct success rate on Alfworld (AW) and Webshop (WS) using available gpt-3.5 models.
@@ 4.5 Metrics
In terms of metrics we use the pre-defined metrics of Alfworld and Webshop, namely success rate (SR). Success is a binary metric per each environment in the respective test sets (135 and 500 respectively). Success in Alfworld means the agent has successfully complete the whole task. In Webshop it means the agent has bought an item that has a hundred percent match with the desired item based on a partially hidden list of attributes of the shopping item (e.g. the colour, size, price, etc.).
@@ 5.1 Alfworld
For Alfworld we present the results for ReAct, AdaPlanner with and without code execution and StateAct (ours), which consists of goal + state + thought + action. We also show StateAct without each of the components (i.e. without goal, state and thought). Interestingly we find, contrary to previous findings, that ‘thought’ or ‘reasoning’ actually sometimes harms the performance.
In Table 2, we can see that StateAct with all goal+state+thought and the correction performs the best. It outperforms ReAct with correction by around 13 points (using the same GPT model for ReAct) and by around 9 points (using the better model for ReAct). StateAct also outperforms ReAct by 22 points when corrections are not used. Furthermore, StateAct even outperforms AdaPlanner by 2.48 points, an approach that uses regex for command mapping (similar to our correction) and code-execution.
Perhaps the most surprising finding is that the simple correction described in Section 4.1.1 leads to a 16 and 23 point jump for ReAct and a 27 point jump for StateAct. This indicates that the model generally performs very well, however, struggles with minute differences in domain specific syntax.
@@
Table 2: Success Rate (SR) on the 135 test-set examples from Alfworld. *gpt-3.5-0301 was deprecated in June 2024. **No-Exec means AdaPlanner without code execution. ***joined means that thought + action are produced at every turn. AdaPlanner results are from Sun et al. (2023). All other results are ‘single run’.
@@ 5.2 Webshop
For Webshop we present results for ReAct and StateAct (ours). Similarly, to Alfworld we also present the results of StateAct without each of goal, state and thought. See Table 3. Interestingly, we see that removing thought produces the highest results and outperforms ReAct by 10 points. Our hypothesis is that domain specific syntax, which is more prevalent in Webshop than Alfworld, conflicts with using verbose thoughts.
@@ Results
Table 3: Success Rate (SR) on the 500 test-set examples from Webshop. *results taken from Yao et al. (2023a). Results are ‘single run’, except for StateAct without thought and ReAct, where we ran the experiment twice and reported the average.
@@ 5.3 Results using additional models
We validate our results further by comparing on additional LLMs. Again we see a significant boost in performance, see Table 4. We ran Gpt-4o-mini and Mixtral [^5] The concrete version of Mixtral was ‘mixtral-8x22b-instruct-v0.1’. (Jiang et al., 2024) on Alfworld. ReAct achieves 68.15 and 72.59 respectively, while StateAct achieves 71.85(+3.7) and 83.70(+11.2).
@@ Results
Table 4: Success Rate (SR) on the 135 test-set examples from Alfworld. Results are ‘single run’. *=The concrete version of Mixtralis ‘mixtral-8x22b-instruct-v0.1’.
@@ 5.4 Summary of results
In conclusion we found that our simple goal-reminding and state tracking approach that purely relies on in-context learning outperforms previous in-context learning approaches by more almost 10 points and even outperforms leading approaches that rely on code-execution. Interestingly, we found that the approaches are quite sensitive to domain specific syntax and that when this is the case ‘thoughts’ that are verbose can harm performance.
@@ 6 Analysis and Ablations
In the results section we discovered that our methods perform better than previous state-of-the-art. This answers the question that we can perform better with in-context learning without resorting to additional tools, data or bigger models. In this section we want to analyse our results further and particularly also answer if our second hypothesis that goal ‘reminding’ and state tracking help with long-range reasoning actually holds. For all ablation studies we focus on Alfworld as it has two favourable properties over Webshop. Firstly, Alfworld has a longer time horizon (50 steps vs. 15 in Webshop), with tasks taking an average of less than 10 steps in Webshop and around 20 to 30 steps in Alfworld. Secondly, Alfworld has much less domain specific syntax and is purely text based, while Webshop has a more specific syntax to follow.
6.1 Do goal reminders help with long range tasks?
For this purpose we compare the original ReAct (thought+action) with just adding the goal in, i.e. StateAct (goal + thought + action). In Figure 5 we can see that while the performance of both ReAct and StateAct goes down as there are more number of steps the goal tracking has a significantly better relative performance as the number of steps increase.
To verify that this actually means that goal tracking helps with performance, as opposed to just increasing the number of steps it takes to solve a task, we calculate the average number of steps for ReAct (ignoring empty ‘thought’ turns, as otherwise ReAct would have even more steps) and StateAct. Table 5 clearly show that ReAct with an average of 38.84 steps to solve an environment is the least efficient and StateAct with an average of 28.96 steps to solve an environment is the most efficient. This shows that not only does goal tracking help with longer range tasks, it also helped with efficiency by shortening the tasks.
Figure 5: Goal vs. No Goal, on the 135 test examples from Alfworld, using gpt-3.5-turbo-1106 without correction.
@@ 6.2 What effects does state-tracking have?
We also analyse whether state tracking helps with long-range reasoning and efficiency. We compared the full StateAct against StateAct without state-tracking as well as ReAct (thought + action) against StateAct with state-tracking added (state + thought + action). In Figure 6 we see that state tracking also helps with long-range reasoning. In fact, we can see that reasoning alone is unable to solve tasks longer than 40 steps, while with state tracking even longer-range tasks can be solved than with goal-tracking alone. Also, looking at Table 5 we see that state-tracking makes the model the most efficient [^6] In terms of cost we found that despite our method using a twice longer prompt, our cost remains similar at around \$8 for the full Alfworld run, since we solve tasks more efficiently and use fewer number of steps..
Therefore we find that explicit state-tracking even further helps with long-range tasks and helps the agent solve the tasks more efficiently than without.
Figure 6: State vs. No State, on the 135 test examples from Alfworld, using gpt-3.5-turbo-1106 with correction.
@@
Table 5: Average number of Steps (Avg. Steps) [lower is better] on the 135 test-set examples from Alfworld. gpt-3.5-1106 for all methods.
6.3 Does the model perform actual state tracking?
We ask ourselves the question if the model is actually performing state-tracking. For that purpose we look at Alfworld and construct a self-verification algorithm that is able to track the state heuristically [^7] On our few-shot prompts it achieves 100% correctness. based on the actions the agent takes. For example if the agent produces the action go to fridge 1 and the environment accepts this action we update the state with current location: fridge 1. We compare the ‘gold’ state against the predicted state. Figure 7 shows that StateAct in fact does correct state-tracking 88% of the time. We also observe that thoughts and goals help the state tracking.
Figure 7: State tracking accuracy for StateAct on 135 test examples of Alfworld using gpt-3.5-1106.
@@ 6.4 Does json structure help with performance?
Since we found that domain specific syntax harms performance, we wondered whether adding a structured format like json would help. For this purpose, we re-ran StateAct on Alfworld, but translated the state into a json format, see Appendix D for more details. Surprisingly, we found that the json format harms performance significantly, see Table 6. However, we also see that corrections help the the json format less, indicating that json helps with syntax, but harms performance.
@@
Table 6: Success Rate (SR) on the 135 test examples from Alfworld. Showing no-json vs. json, gpt-3.5-1106.
@@ 7 Conclusion and future work
We propose a novel method StateAct, using our ‘chain-of-states’, based on in-context learning alone and establish a new state-of-the-art for agents that do not perform training, even against methods that use code-execution. The method outperforms the previous state-of-the-art, that uses in-context learning alone, between 9% and 20% given different models and tasks and outperform in-context learning with tools (code-execution) by 3%. We also show that explicit state-tracking and goal reminders make the model more efficient as well as significantly help with longer range tasks.
We found that ‘thoughts’ or explicit reasoning do not always help performance. It would be very interesting to systematise ‘thought’ and ‘states’ and understand what contributes positively and why. Also, inspired by the good results of StateAct, it is interesting to see what other improvements can be done without resorting to training, larger model or external tools. Finally, problems related to domain specific syntax are also an interesting avenue of future work.
@@ 8.2 Hallucinations in LLMs
As LLM-based agents become more powerful and therefore more pervasive in our daily lives ‘hallucinations’ of LLMs can be very harmful Wei et al. (2024). We hope that explicit state-tracking presented in this work can also lead to future work that can reduce ‘hallucinations.’
@@ 9.1 Languages and evaluation benchmarks
9.2 Reasoning traces rely on human judgement
Our prompts require human annotations, as such there is a natural bias present. This can have both task-performance implications as well as ethical implications.
@@ A.1 Environment Types
The ‘clean’ task, e.g. Task: Put a clean apple on table, requires the agent to first find the apple, then clean it (in the sinkbasin) and then put it on a table.
The ‘heat’ task, e.g. Task: Put a hot pie on table, requires the agent to first find the pie, then heat it (on the stoveburner) and then put it on a table.
@@ A.3 License
Alfworld has the permissible MIT license, we used it in line with the license.
@@ B.2 Prodcuts and attributes
Webshop has over 1 million real-world products across 5 main categories (fashion, makeup, electronics, furniture, and food) and 113 sub-categories.
@@ B.3 License
Webshop has the permissible Princeton license, we used it in line with the license.
@@
A prompt is given in Appendix E.
Appendix D StateAct Json Format
We translate the text based StateAct prompt:
@@ Appendix D StateAct Json Format
Appendix E StateAct Prompts
E.1 Example Alfworld prompt for the ‘heat’ task
@@
E.2 Example Webshop prompt

### v2

@@ 4.4 Models
In line with previous work, we focus our main experiments are based on the API-based LLMs to compare performance fairly. Concretely, ‘gpt-3.5’ level models were used in previous work including current state-of-the-art ActRe (Yang et al., 2024b) and state-of-the-art for in-context learning ReAct (Yao et al., 2023b). Many OpenAI models have become deprecated; notably, all models from ReAct and AdaPlanner Sun et al. (2023) davinci-002, gpt-3.5-turbo-0301, and gpt-3.5-turbo-0613 are now deprecated. Therefore, we re-implemented ReAct and ran the experiments to determine the most suitable model, see Table 1. We establish that gpt-3.5-turbo-1106 is the best performing (from the ones that remain available) on ReAct and we therefore chose this model for our main experiments.
In our work, we also compare our method using newer state-of-the-art models to show that our method generalises to new models. Specificially, we use the gpt-4o-mini [^4] Again, we chose the smaller model due to cost. and Mixtral (Jiang et al., 2024) models. We note that full GPT-4 level models are prohibitively expensive [^5] A single evaluation run on Alfworld costs approx. \$8 using gpt-3.5 and ReAct, whilst gpt-4 costs 10+ times more.. We use temperature 0 for all experiments and sample only the top 1 response, see Appendix C for the exact settings.
@@
Table 1: ReAct success rate on Alfworld (AW) and Webshop (WS), using available gpt-3.5 models.
@@ 4.5 Metrics
In terms of metrics, we use the pre-defined metrics of Alfworld and Webshop, namely success rate (SR). Success is a binary metric per each environment in the respective test sets (134 for Alfworld and 500 for Webshop respectively). Success in Alfworld means the agent has successfully completed the whole task. In Webshop, it means the agent has bought an item that has a one hundred percent match with the desired item based on a partially hidden list of shopping item attributes (e.g. the colour, size, price, etc.).
@@ 5.1 Alfworld
For Alfworld, we present the results for ReAct, AdaPlanner with and without code execution, and StateAct (ours), which consists of goal + state + thought + action. We also show StateAct without each of the components (i.e. without goal, state and thought). Interestingly, we find that, contrary to previous findings, ‘thought’ or ‘reasoning’ actually sometimes hinders the performance.
In Table 2, we can see that StateAct with all goal + state + thought and the correction performs the best. It outperforms ReAct with correction by around 13 points (when using the same GPT model for ReAct) and by around 9 points (when using the better model for ReAct). StateAct also outperforms ReAct by 22 points when corrections are not used. Furthermore, StateAct even outperforms AdaPlanner by 2.48 points, an approach that uses regex for command-mapping (similar to our correction) along with using code-execution.
Perhaps the most surprising finding is that the simple correction described in Section 4.1.1 leads to a 16- and 23-point jump for ReAct and a 27-point jump for StateAct. This indicates that the model generally performs very well; however, it struggles with minute differences in domain-specific syntax.
@@
Table 2: Success Rate (SR) on the 134 test-set examples from Alfworld. *gpt-3.5-0301 was deprecated in June 2024. **No-Exec means AdaPlanner without code execution. ***joined means that thought + action are produced at every turn. AdaPlanner results are from Sun et al. (2023). All other results are single run.
@@ 5.2 Webshop
For Webshop, we present the results for ReAct and StateAct. Similarly to Alfworld, we also present the results of StateAct without each of goal, state and thought, see Table 3. Interestingly, we see that removing thought produces the highest results and outperforms ReAct by 10 points. Our hypothesis is that domain-specific syntax, which is more prevalent in Webshop than Alfworld, conflicts with using verbose thoughts.
@@ Results
Table 3: Success Rate (SR) on the 500 test-set examples from Webshop. *results taken from Yao et al. (2023a). For StateAct without thought and ReAct Results are the average of two runs. Other Results are single run.
@@ 5.3 Results using additional models
We validate our results further with additional LLMs. Again we see a significant boost in performance, see Table 4. We experimented with gpt-4o-mini and Mixtral [^6] The concrete version of Mixtral was ‘mixtral-8x22b-instruct-v0.1’. (Jiang et al., 2024) on Alfworld. ReAct achieves 68.15 and 72.59 respectively, while StateAct achieves 71.85(+3.7) and 83.70(+11.2).
@@ Results
Table 4: Success Rate (SR) on the 134 test-set examples from Alfworld. Results are ‘single run’. *=The concrete version of Mixtral is ‘mixtral-8x22b-instruct-v0.1’.
@@ 5.4 Summary of results
In conclusion, we found that our simple goal-reminding and state-tracking approach that purely relies on in-context learning outperforms previous in-context learning approaches by more than 10 points and even outperforms leading approaches that rely on code-execution. Interestingly, we found that the approaches are quite sensitive to domain-specific syntax and, when this is the case, ‘thoughts’ that are verbose can hinder performance.
@@ 6 Analysis and Ablations
In the results section, we discovered that our methods perform better than previous state-of-the-art. Our method shows strong performance with in-context learning without resorting to additional tools, data or bigger models. In this section, we analyse our results further and also answer our second hypothesis that goal ‘reminding’ and state-tracking help with long-range reasoning. For all ablation studies, we focus on Alfworld as it has two favourable properties over Webshop. Firstly, Alfworld has a longer time horizon (50 steps vs. 15 in Webshop), with tasks taking, on average, less than 10 steps in Webshop and around 20 to 30 steps in Alfworld. Secondly, Alfworld has much less domain-specific syntax and is purely text-based, while Webshop has a more specific syntax to follow.
6.1 Do goal reminders help with long-range tasks?
For this purpose, we compare the original ReAct (thought + action) with the goal included, i.e. StateAct (goal + thought + action). In Figure 5 we can see that, while the performance of both ReAct and StateAct goes down as there are more steps, the goal tracking has a substantially better relative performance as the number of steps increases. This finding is in-line with our original motivation that LLM agents deteriorate in performance as the prompts and interactions get longer.
To verify that this actually means that goal tracking helps with performance, as opposed to just increasing the number of steps it takes to solve a task, we calculate the average number of steps for ReAct [^7] We ignore ‘thought’ turns for ReAct as otherwise ReAct would have even more steps. and StateAct. Table 5 clearly show that ReAct, with an average of 38.84 steps to solve an environment, is the least efficient whilst StateAct, with an average of 28.96 steps to solve an environment, is the most efficient. This shows that not only does goal reminders help with longer range tasks, it also helps with efficiency, by shortening the tasks. See Appendix E for more discussion.
Figure 5: Goal vs. No Goal, on the test set of Alfworld, using gpt-3.5-turbo-1106 without correction.
@@ 6.2 What effects does state-tracking have?
We also analyse whether state-tracking helps with long-range reasoning and efficiency. We compare the full StateAct against StateAct without state-tracking, as well as comparing ReAct (thought + action) against StateAct with state-tracking (state + thought + action). In Figure 6 we see that state-tracking also helps with long-range reasoning. In fact, we can see that reasoning alone is unable to solve tasks longer than 40 steps, while with both state tracking and goal-tracking longer-range tasks can be solved. Also, looking at Table 5 we see that state-tracking makes the model the most efficient [^8] In terms of cost, we found that, despite our method using a twice-longer prompt, our cost remains similar, at around \$8 for the full Alfworld run, since we solve tasks more efficiently and use fewer steps.. Concretely, ReAct has 24/134 examples that are in the bucket ‘40-50’ and solves 0, while our method has 6/134 examples and solves 4.
Therefore, we find that explicit state-tracking helps with long-range tasks and enables the agent to solve the tasks more efficiently.
Figure 6: State vs. No State, on the 134 test examples from Alfworld. gpt-3.5-turbo-1106 with correction.
@@
Table 5: Average number of Steps (Avg. Steps) [lower is better] on test-set of Alfworld, using gpt-3.5-1106.
6.3 Does the model perform actual state-tracking?
We ask ourselves whether the model is actually performing state-tracking. For that purpose, we look at Alfworld and construct a self-verification algorithm that is able to track the state heuristically [^9] See Appendix D for details. based on the actions the agent takes. For example, if the agent produces the action go to fridge 1 and the environment accepts this action, we update the state with current location: fridge 1. We compare the ‘gold’ state against the predicted state. Figure 7 shows that StateAct does in fact correct state-tracking 88% of the time. We also observe that thoughts and goals help the state-tracking.
Figure 7: State-tracking accuracy for StateAct on 134 test examples of Alfworld, using gpt-3.5-1106.
@@ 6.4 Does json structure help with performance?
Since we found that domain-specific syntax hinders performance, we investigated whether adding a structured format like json would help. For this purpose, we re-ran StateAct on Alfworld, but translated the state into a json format, see Appendix G for more details. Surprisingly, we found that the json format hinders performance significantly, see Table 6. However, we also see that corrections help the json format less, indicating that json helps with syntax but hinders performance.
@@
Table 6: No-json vs. json. Success Rate (SR) on the test set of Alfworld, using gpt-3.5-1106.
@@ 7 Conclusion and future work
Inspired by the challenge that LLMs and LLM agents struggle with long context, we propose a novel in context learning method for LLM agents - StateAct.
We introduce ‘chain-of-states’, which relies on in-context learning alone and establish a new state-of-the-art for agents that do not perform training, even against methods that use code-execution. StateAct outperforms the previous state-of-the-art that uses in-context learning alone, by between 9% and 20% given different models and tasks, and outperforms in-context learning with tools (i.e. code-execution) by 3%, while performing on par with methods that require additional training data.
In conclusion, we show that explicit state-tracking and goal reminders help the model solve long range reasoning tasks and interestingly also make the model more efficient.
@@ 8.2 Hallucinations in LLMs
As LLM-based agents become more powerful and therefore more pervasive in our daily lives, ‘hallucinations’ in LLMs can be very harmful Wei et al. (2024). We hope that explicit state-tracking presented in this work can also lead to future work to reduce ‘hallucinations.’
@@ 9.1 Languages and evaluation benchmarks
9.2 Reasoning traces that rely on human judgement
Our prompts require human annotations; as such, there is a natural bias present. This can have both task-performance implications as well as ethical implications.
@@ A.1 Environment Types
The ‘clean’ task, e.g. Task: Put a clean apple on table, requires the agent to first find the apple, then clean it (in the sink/basin) and then put it on a table.
The ‘heat’ task, e.g. Task: Put a hot pie on table, requires the agent to first find the pie, then heat it (on the stove/burner) and then put it on a table.
@@ A.3 License
Alfworld has the permissible MIT license; we used it in line with the license.
@@ B.2 Prodcuts and attributes
Webshop has over 1 million real-world products across 5 main categories (fashion, makeup, electronics, furniture and food) and 113 sub-categories.
@@ B.3 License
Webshop has the permissible Princeton license; we used it in line with the license.
@@
A prompt is given in Appendix H.
Appendix D Heuristic State-tracking Explained
Heuristic state tracking is based on the idea that the state can be inferred automatically if one follows the actions of the agent and observations of the environments. Specifically, the state at time $t$ for StateAct depends on the state at time t-1 and the action $a_{t-1}$ and observation $o_{t}$. For example, if the state at time $t$ the ‘current location’ of the state is set to table 1 and the action is go to fridge 1 and the observation is successful, then the ‘current location’ can be updated to be fridge 1 automatically. This rule based ‘state-tracking’ is how the heuristics work.
Appendix E ‘Step Length Analysis’ Discussion
An alternative to calculating and comparing step length could be ‘gold solutions’ to measure optimal step length and optimality of an agent. We see two issues. Firstly, the annotation cost of creating gold solutions. Secondly, it is not clear what the gold solution should be. Concretely in Alfworld, an ‘oracle’ solution could have very few steps as it would immediately go to the location of the ‘hidden’ object, while a ‘non-oracle’ expert solution would have more steps as more locations would be searched. Thus a ‘difficulty’ measure would be needed instead, but it is ambiguous.
Appendix F Potential Future Work Directions
We found that ‘thoughts’ or explicit reasoning do not always help performance. It would be very interesting to systematise ‘thought’ and ‘states’ and to understand what contributes positively and the reasons why. Also, inspired by the positive results of StateAct, it is interesting to see what other improvements can be made without resorting to training, larger models or external tools. Finally, problems related to domain-specific syntax are also an interesting avenue for future work.
Appendix G StateAct Json Format
We translate the text-based StateAct prompt:
@@ Appendix G StateAct Json Format
Appendix H StateAct Prompts
H.1 Example Alfworld prompt for the ‘heat’ task
@@
H.2 Example Webshop prompt